# Gaze-guided Hand-Object Interaction Synthesis: Dataset and Method

### v4

@@ Abstract
Gaze plays a crucial role in revealing human attention and intention, particularly in hand-object interaction scenarios, where it guides and synchronizes complex tasks that require precise coordination between the brain, hand, and object. Motivated by this, we introduce a novel task: Gaze-Guided Hand-Object Interaction Synthesis, with potential applications in augmented reality, virtual reality, and assistive technologies. To support this task, we present GazeHOI, the first dataset to capture simultaneous 3D modeling of gaze, hand, and object interactions. This task poses significant challenges due to the inherent sparsity and noise in gaze data, as well as the need for high consistency and physical plausibility in generating hand and object motions. To tackle these issues, we propose a stacked gaze-guided hand-object interaction diffusion model, named GHO-Diffusion. The stacked design effectively reduces the complexity of motion generation. We also introduce HOI-Manifold Guidance during the sampling stage of GHO-Diffusion, enabling fine-grained control over generated motions while maintaining the data manifold. Additionally, we propose a spatial-temporal gaze feature encoding for the diffusion condition and select diffusion results based on consistency scores between gaze-contact maps and gaze-interaction trajectories. Extensive experiments highlight the effectiveness of our method and the unique contributions of our dataset.
Introduction
Gaze serves as a significant behavioral signal, directly reflecting human attention distribution and cognitive processes (Schneider 1995; Gottlieb et al. 2013; Desanghere and Marotta 2015). Extensive research has been conducted on gaze, including gaze estimation (Abdelrahman et al. 2023; Guan et al. 2023; Kothari et al. 2021; Fischer, Chang, and Demiris 2018) and human action prediction based on gaze (Kratzer et al. 2020; Zheng et al. 2022; DelPreto et al. 2022; Ragusa et al. 2021; Eren Akbiyik et al. 2023). While these studies have explored gaze in various applications, they have primarily focused on broad human motions, with limited attention given to hand movements and detailed dynamic interactions within fine-grained environments. Furthermore, gaze is particularly crucial in hand-object interactions, as it reveals ”when, where, and how” the user interacts with the object. Figure LABEL:fig:teaser(a) illustrates how gaze supports long-horizon synthesis for complex tasks, while Figure LABEL:fig:teaser(b) shows that gaze indicates how humans interact with objects using different parts and interaction types. By effectively identifying objects and areas of focus, gaze enables the seamless integration of human intentions and motions, ensuring precise coordination and consistency between the brain, hands, and objects. This coordination is essential for executing complex interactions, such as grasping, manipulating, or positioning objects, thereby enhancing both the efficiency and accuracy of these motions.
@@ Introduction
Existing hand-object datasets are limited to interactions between hands and objects, often neglecting gaze information. To fill this gap, we introduce GazeHOI, the first dataset that simultaneously integrates 3D modeling of gaze, hand, and object interactions. GazeHOI presents significant challenges due to the diverse shapes and sizes of objects, ranging from small items like chess pieces to larger, assemblable furniture. Additionally, the dataset covers a wide range of complex tasks, including repositioning objects, selecting specific targets within cluttered environments, organizing disordered items, and assembling furniture.
@@
Figure 1: Automatic data processing pipeline. (a) displays 12-view images from the raw video. (b) uses mediapipe(Lugaresi et al. 2019) to get hand 2D joints and triangulation to obtain the 3D joints. (c) shows the ego view with a gaze point. (d) illustrates the objects acquired by the scanner and the marker tracking process. (e) shows the result of hand-object motion.
Related Work
Gaze for Visual Perception
Gaze is fundamental to reveal intentional cues and offer insights into attention, interest, and cognitive processes. Due to the expense and accuracy limitations of gaze tracking equipment, the previous researchers have primarily explored methods for estimating gaze from a third-person perspective images or videos (Abdelrahman et al. 2023; Guan et al. 2023; Kothari et al. 2021; Fischer, Chang, and Demiris 2018). With the recent advancements in gaze tracking technology, acquiring accurate gaze information has become relatively more convenient, leading to the emergence of numerous datasets (Kratzer et al. 2020; Zheng et al. 2022; DelPreto et al. 2022; Ragusa et al. 2021; Eren Akbiyik et al. 2023) containing gaze annotations. Researchers have moved beyond solely acquiring gaze data and are now exploring methods for extracting rich information inherent in gaze, such as inferring the ego-trajectory of a driver’s vehicle (Eren Akbiyik et al. 2023) or predicting the interaction action labels (Liu et al. 2020). MoGaze and GIMO(Kratzer et al. 2020; Zheng et al. 2022) propose leveraging gaze to predict human motions within the scene, which reveals the potential of gaze in virtual reality (VR) and human-robot interaction scenarios. They solely focus on human motions, but fine-grained hand interaction generation is crucial for the realism of VR experiences. Therefore, in this work, we introduce the task of gaze-guided hand-object interaction motion synthesis.
Hand-Object Interaction Dataset
@@
Numerous datasets (Hampali et al. 2020; Chao et al. 2021; Kwon et al. 2021; Yang et al. 2022; Liu et al. 2022; Fan et al. 2023; Liu et al. 2024; Zhang et al. 2023; Zhao et al. 2023; Zhan et al. 2024) dedicated to hand-object interactions have been developed to facilitate the capture of hand and object poses and to deepen our understanding of these interactions, as shown in Table 1. Most datasets only capture simple interaction patterns,like grasping and moving, which is common but not comprehensive enough. HOI4D (Liu et al. 2022) and ARCTIC (Fan et al. 2023) explore the interaction of articulated objects. TACO (Liu et al. 2024) and Oakink2 (Zhan et al. 2024) captures scenes of simultaneous interactions with multiple objects. Furthermore, existing datasets focus exclusively on the relationship between hands and objects, neglecting the role of gaze. To address this gap, we introduce a comprehensive egocentric dataset enriched with detailed gaze annotations. This dataset captures the dynamics of hand-object interactions across a wide range of tasks, from arranging chess pieces to assembling furniture — scenarios not covered in existing datasets. Our goal is to bridge the shortcomings of current datasets by providing a resource that reflects the complexity of real-world interactions and emphasizes the importance of gaze in predicting human intent and managing multi-object manipulation.
@@
HO-3D (Hampali et al. 2020) $\times$ $\times$ $\times$ $\times$ 1-4 78K
DexYCB (Chao et al. 2021) $\times$ $\times$ $\times$ $\times$ 8 582K
H2O (Kwon et al. 2021) $\times$ $\checkmark$ $\times$ $\times$ 5 571K
OakInk (Yang et al. 2022) $\times$ $\times$ $\times$ $\times$ 4 230K
HOI4D (Liu et al. 2022) $\times$ $\checkmark$ $\times$ $\times$ 2 2.4M
ARCTIC (Fan et al. 2023) $\times$ $\checkmark$ $\times$ $\times$ 9 2.1M
TACO (Liu et al. 2024) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 13 5.2M
OAKINK2 (Zhan et al. 2024) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 4 4.01M
@@
Table 1: Comparison of GazeHOI with existing 4D hand-object interaction datasets.
Hand-Object Motion Synthesis
In the field of hand-object motion synthesis, prior efforts have mainly centered on frame-wise generation (Li et al. 2022; Jiang et al. 2021; Liu et al. 2023; Grady et al. 2021). For dynamic hand-object interaction, previous methods typically rely on motion references, such as trajectories (Zhang et al. 2021; Karunratanakul et al. 2023; Taheri et al. 2024; Zhou et al. 2024; Li, Wu, and Liu 2023; Zheng et al. 2023) pr grasp references (Christen et al. 2022), or only focus on approaching stage (Taheri et al. 2022; Ghosh et al. 2023). Recently, there are several works that study the hand-object interaction generation directly based on text (Cha et al. 2024; Christen et al. 2024), significantly reducing the reliance on specific conditions. Building on advancements in 3D gaze acquisition technology, our work further relaxes these constraints by generating hand-object interactions directly from gaze data. Compared to text, gaze-based generation aligns more closely with the user’s intent, resulting in more intuitive hand-object interactions.
GazeHOI Dataset
@@
Data Collection
@@
Data Annotation
3D Hand Pose.
12 synchronized ZCAM cameras afford us an exhaustive 3D perspective of the hand, as shown in Figure 1(a). By harnessing the capabilities of MediaPipe (Lugaresi et al. 2019), we execute 2D hand keypoint tracking across individual video frames. Through the application of triangulation algorithms, we transition the 2D keypoints observed from disparate views to their respective 3D coordinates in global space shown in Figure 1(b). Subsequently, we derive the corresponding MANO (Romero, Tzionas, and Black 2017) parameters through an optimization process that accounts for joint alignment, realistic anatomical constraints.
Object 6D Pose.
@@
Problem Definition
We define a gaze-consistent hand-object interaction sequence of length $l$ as $\mathbf{S}=\langle\mathbf{G},\mathbf{H},\mathbf{O}\rangle$, where $\mathbf{G}=\{\mathbf{g}_{i}\}_{i=1}^{l}$ represents the gaze sequence, with each $\mathbf{g}_{i}\in\mathbb{R}^{3}$ denoting the gaze point in the $i$-th frame. $\mathbf{H}=\{\mathbf{h}_{l}^{i},\mathbf{h}_{r}^{i}\}_{i=1}^{l}$ denotes the hand motion sequences, comprising left-hand motions $\mathbf{h}_{l}^{i}$ and right-hand motions $\mathbf{h}_{r}^{i}$. Each hand motion $\mathbf{h}^{i}=\{\mathbf{R}_{h}^{i},\mathbf{T}_{h}^{i},\bm{\theta}^{i},\bm{\beta}\}\in\mathbb{R}^{61}$ follows the MANO (Romero, Tzionas, and Black 2017) parameterization, where $\mathbf{R}_{h}^{i}$ is the rotation, $\mathbf{T}_{h}^{i}$ is the translation, $\bm{\theta}^{i}$ represents the pose and $\bm{\beta}$ represents shape parameters; $\mathbf{O}=\{\mathbf{o}^{i}\}_{i=1}^{l}$ is the object motion sequence, where each $\mathbf{o}^{i}=\{\mathbf{R}_{o}^{i},\mathbf{T}_{o}^{i}\}\in\mathbb{SE}(3)$ denotes the object’s pose, with $\mathbf{R}_{o}^{i}$ as the rotation and $\mathbf{T}_{o}^{i}$ as the translation. Additionally, we define the object geometry as a set of vertices $\mathbf{P}_{o}=\{\mathbf{p}^{n}_{o}\}_{n=1}^{N}\in\mathbb{R}^{N\times 3}$, where $N$ represents the number of object vertices. Given the input =$\mathcal{I}=\{\mathbf{G},\mathbf{h}^{1}_{l},\mathbf{h}^{1}_{r},\mathbf{o}^{1},\mathbf{P}_{o}\}$, our objective is to derive a subsequent hand-object interaction sequence $\langle\mathbf{H},\mathbf{O}\rangle$ that is not only consistent with the gaze but also exhibits natural and realistic interactions between the hand and the object.
GHO-Diffusion: Gaze-guided Hand-Object Diffusion
@@
The diffusion model contains a fixed forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ which gradually adds Gaussian noise to clean data $\mathbf{x}_{0}$ until it becomes pure noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and train a denoiser $p_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ to gradually denoise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to generate $\mathbf{x}_{0}$ in the reverse process. Instead of predicting $\bm{\epsilon}_{t}$ at each time step $t$, we follow MDM (Tevet et al. 2022) to predict $\hat{\mathbf{x}}_{0}$ and optimizing $p_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ using the simple objective function:
@@
Stage-1: Object Dynamics Synthesis with Spatial-Temporal Gaze Encoding.
The gaze point locations are closely related to the object surface vertices, and since object motions have fewer degrees of freedom compared to hand motions, so we generate the object motions first. Gaze is a sparse and noisy signal, making it challenging to encode effectively. Drawing inspiration from GIMO (Zheng et al. 2022), we design a gaze-guided spatial-temporal feature encoding module for object diffusion. We define $\mathbf{f}^{k}_{o}$ as the $k$-th local per-point object spatial feature extracted by PointNet++ (Qi et al. 2017). Then according to the spatial relationship between gaze and object points, we can calculate gaze spatial feature:
@@
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are trainable weight matrices, and $d$ is the dimension of the feature. We leverage the gaze spatial-temporal features $\mathbf{c}_{g}$, along with the object geometry $\mathbf{P}$ and initial object pose $\mathbf{o}^{1}$ as stage-1 diffusion condition and employ the objective function in Equation 4 as Stage-1 training loss to generate object motions $\hat{\mathbf{O}}$:
@@
where $\mathcal{L}_{\text{simple}}$ is defined by Equation 1, $\mathcal{L}_{\text{trans}}$ indicates object translation loss, while $\mathcal{L}_{\text{verts}}$ represents object vertices loss. The details of these loss terms are provided in the Appendix.
Stage-2: Hand Kinematic Synthesis with HOI Guidance

### v5

@@ Abstract
Gaze plays a crucial role in revealing human attention and intention, particularly in hand-object interaction scenarios, where it guides and synchronizes complex tasks that require precise coordination between the brain, hand, and object. Motivated by this, we introduce a novel task: Gaze-Guided Hand-Object Interaction Synthesis, with potential applications in augmented reality, virtual reality, and assistive technologies. To support this task, we present GazeHOI, the first dataset to capture simultaneous 3D modeling of gaze, hand, and object interactions. This task poses significant challenges due to the inherent sparsity and noise in gaze data, as well as the need for high consistency and physical plausibility in generating hand and object motions. To tackle these issues, we propose a stacked gaze-guided hand-object interaction diffusion model, named GHO-Diffusion. The stacked design effectively reduces the complexity of motion generation. We also introduce HOI-Manifold Guidance during the sampling stage of GHO-Diffusion, enabling fine-grained control over generated motions while maintaining the data manifold. Additionally, we propose a spatial-temporal gaze feature encoding for the diffusion condition and select diffusion results based on consistency scores between gaze-contact maps and gaze-interaction trajectories. Extensive experiments highlight the effectiveness of our method and the unique contributions of our dataset. More details in https://takiee.github.io/gaze-hoi/.
Index Terms:
Hand-object interaction, gaze, dataset, motion generation
I Introduction
Gaze serves as a significant behavioral signal, directly reflecting human attention distribution and cognitive processes [1, 2, 3]. Extensive research has been conducted on gaze, including gaze estimation [4, 5, 6, 7, 8, 9, 10] and human action prediction based on gaze [11, 12, 13, 14, 15]. While these studies have explored gaze in various applications, they have primarily focused on broad human motions, with limited attention given to hand movements and detailed dynamic interactions within fine-grained environments. Furthermore, gaze is particularly crucial in hand-object interactions, as it reveals ”when, where, and how” the user interacts with the object. Figure LABEL:fig:teaser(a) illustrates how gaze supports long-horizon synthesis for complex tasks, while Figure LABEL:fig:teaser(b) shows that gaze indicates how humans interact with objects using different parts and interaction types. By effectively identifying objects and areas of focus, gaze enables the seamless integration of human intentions and motions, ensuring precise coordination and consistency between the brain, hands, and objects. This coordination is essential for executing complex interactions, such as grasping, manipulating, or positioning objects, thereby enhancing both the efficiency and accuracy of these motions.
@@ I Introduction
Existing hand-object datasets[16, 17, 18, 19, 20, 21, 22, 23, 24, 25] are limited to interactions between hands and objects, often neglecting gaze information. To fill this gap, we introduce GazeHOI, the first dataset that simultaneously integrates 3D modeling of gaze, hand, and object interactions. GazeHOI presents significant challenges due to the diverse shapes and sizes of objects, ranging from small items like chess pieces to larger, assemblable furniture. Additionally, the dataset covers a wide range of complex tasks, including repositioning objects, selecting specific targets within cluttered environments, organizing disordered items, and assembling furniture.
@@
Figure 1: Automatic data processing pipeline. (a) displays 12-view images from the raw video. (b) uses mediapipe[26] to get hand 2D joints and triangulation to obtain the 3D joints. (c) shows the ego view with a gaze point. (d) illustrates the objects acquired by the scanner and the marker tracking process. (e) shows the result of hand-object motion.
II Related Work
II-A Gaze for Perception
Gaze serves as a key behavioral signal, revealing intentional cues and providing insights into attention, interest, and cognitive processes. Early research, constrained by the expense and accuracy limitations of gaze tracking equipment, focused on estimating gaze from third-person perspective images or videos [4, 8, 9, 10, 5, 6, 7] to predict human attention regions. With advancements in gaze tracking technology[27], acquiring accurate gaze information has become more accessible, leading to the emergence of diverse datasets [11, 12, 13, 14, 15, 28, 29, 30] annotated with gaze information. Recent studies have expanded from simply capturing gaze to leveraging its contextual richness, such as inferring vehicle ego-trajectory [15] or predicting interaction action labels [31]. However, existing gaze-motion datasets [11, 12, 13, 28, 29, 30] and methods largely overlook the crucial relationship between gaze and hand-object interaction—key for fine-grained intent inference and enhancing realism in applications like VR. From a dataset perspective, current datasets either lack hand pose capture [11, 29] or object pose capture [12, 28, 13, 30] as shown in Table I. From a task perspective, most focus on human motion prediction[32, 33, 12], neglecting the precision required for hand-object interactions, such as realistic contact, penetration avoidance, and plausible physical dynamics. Existing human motion prediction methods also rely heavily on historical motion sequences, using gaze data as auxiliary input within gaze-motion fusion encoders. These approaches fail to fully exploit the spatial relationships between gaze and objects and the temporal dynamics of gaze in hand-object interactions. To address these gaps, we present the first dataset to simultaneously capture 3D gaze, hand, and object interactions, with the novel task of gaze-guided hand-object interaction motion synthesis from an initial state, without relying on historical motions. To effectively capture the intentionality inherent in gaze, we design a gaze-guided spatial-temporal feature encoding as diffusion condition and develop a mechanism to select the diffusion results based on consistency scores computed between gaze-contact maps and gaze-interaction trajectories.
TABLE I: comparison of GazeHOI with Existing gaze datasets
Gaze Object pose Hand pose
MoGaze[11] $\checkmark$ $\checkmark$ $\times$
GIMO[12] $\checkmark$ $\times$ $\checkmark$
ActionSense[13] $\checkmark$ $\times$ $\checkmark$
MECCANO[14] $\checkmark$ $\times$ $\times$
HoloAssist[28] $\checkmark$ $\times$ $\checkmark$
ADT[29] $\checkmark$ $\checkmark$ $\times$
EgoBody[30] $\checkmark$ $\times$ $\checkmark$
GazeHOI $\checkmark$ $\checkmark$ $\checkmark$
II-B Hand-Object Interaction Dataset
@@
Numerous datasets [16, 17, 18, 19, 20, 21, 22, 23, 24, 25] dedicated to hand-object interactions have been developed to facilitate the capture of hand and object poses and to deepen our understanding of these interactions, as shown in Table II. Most datasets only capture simple interaction patterns,like grasping and moving, which is common but not comprehensive enough. HOI4D [20] and ARCTIC [21] explore the interaction of articulated objects. TACO [22] and OakInk2 [25] captures scenes of simultaneous interactions with multiple objects. Furthermore, existing datasets focus exclusively on the relationship between hands and objects, neglecting the role of gaze. To address this gap, we introduce a comprehensive egocentric dataset enriched with detailed gaze annotations. This dataset captures the dynamics of hand-object interactions across a wide range of tasks, from arranging chess pieces to assembling furniture — scenarios not covered in existing datasets. Our goal is to bridge the shortcomings of current datasets by providing a resource that reflects the complexity of real-world interactions and emphasizes the importance of gaze in predicting human intent and managing multi-object manipulation.
TABLE II: Comparison of GazeHOI with existing 4D hand-object interaction datasets.
@@
HO-3D [16] $\times$ $\times$ $\times$ $\times$ 1-4 78K
DexYCB [17] $\times$ $\times$ $\times$ $\times$ 8 582K
H2O [18] $\times$ $\checkmark$ $\times$ $\times$ 5 571K
OakInk [19] $\times$ $\times$ $\times$ $\times$ 4 230K
HOI4D [20] $\times$ $\checkmark$ $\times$ $\times$ 2 2.4M
ARCTIC [21] $\times$ $\checkmark$ $\times$ $\times$ 9 2.1M
TACO [22] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 13 5.2M
OAKINK2 [25] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 4 4.01M
@@
II-C Hand-Object Motion Synthesis
In the field of hand-object motion synthesis, prior efforts have mainly centered on frame-wise generation [34, 35, 36, 37]. For dynamic hand-object interaction, previous methods typically rely on motion references, such as trajectories [38, 39, 40, 41, 42, 43] pr grasp references [44], or only focus on approaching stage [45, 46]. Recently, there are several works that study the hand-object interaction generation directly based on text [47, 48], significantly reducing the reliance on specific conditions. Building on advancements in 3D gaze acquisition technology, our work further relaxes these constraints by generating hand-object interactions directly from gaze data. Compared to text, gaze-based generation aligns more closely with the user’s intent, resulting in more intuitive hand-object interactions.
III GazeHOI Dataset
@@
III-A Dataset Hardware Setup
To construct the GazeHOI dataset, we developed a data collection platform encompassing a 1.6m $\times$ 1.2m area, equipped with a multi-camera system. This platform comprises four distinct equipment systems. The ZCAM system includes 12 ZCAM E2 cameras that have been synchronized through hardware, capturing RGB signals at a resolution of 3840 x 2160 at 30fps, which are utilized for subsequent hand pose acquisition. The Optitrack system consists of 8 hardware-synchronized Optitrack px13W cameras, which use infrared to capture the positions of reflective markers at a frame rate of 120fps, facilitating the tracking of object poses. The Scanner system, comprising two EinScan Pro 2X and EinScan SE scanners, is dedicated to collecting geometric information about objects. The Ego system, primarily aimed at gathering gaze data, employs a Pupil Core connected to a RealSense D455, simultaneously recording gaze information along with RGBD data from an ego-centric perspective. We devise a sophisticated workflow to achieve coordinate system registration among the 4 equipment systems. Each system is tasked with collecting specific types of data, and we have devised a sophisticated workflow to achieve coordinate system registration among the 4 equipment systems through reflective markers. We attached eight 10mm diameter markers on the surface of the RealSense and scanned it to obtain its 3D model. By annotating the position of each marker in every RGB viewpoint and triangulating these positions, we could ascertain the location of markers within each system, thereby aligning the coordinate systems of multiple devices.
III-B Data Collection
@@
III-C Data Annotation
III-C1 3D Hand Pose
12 synchronized ZCAM cameras afford us an exhaustive 3D perspective of the hand, as shown in Figure 1(a). By harnessing the capabilities of MediaPipe [26], we execute 2D hand keypoint tracking across individual video frames. Through the application of triangulation algorithms, we transition the 2D keypoints observed from disparate views to their respective 3D coordinates in global space shown in Figure 1(b). Subsequently, we derive the corresponding MANO [49] parameters through an optimization process that accounts for joint alignment, realistic anatomical constraints.
III-C2 Object 6D Pose
@@ IV Method
IV-A Problem Definition
We define a gaze-consistent hand-object interaction sequence of length $l$ as $\mathbf{S}=\langle\mathbf{G},\mathbf{H},\mathbf{O}\rangle$, where $\mathbf{G}=\{\mathbf{g}_{i}\}_{i=1}^{l}$ represents the gaze sequence, with each $\mathbf{g}_{i}\in\mathbb{R}^{3}$ denoting the gaze point in the $i$-th frame. $\mathbf{H}=\{\mathbf{h}_{l}^{i},\mathbf{h}_{r}^{i}\}_{i=1}^{l}$ denotes the hand motion sequences, comprising left-hand motions $\mathbf{h}_{l}^{i}$ and right-hand motions $\mathbf{h}_{r}^{i}$. Each hand motion $\mathbf{h}^{i}=\{\mathbf{R}_{h}^{i},\mathbf{T}_{h}^{i},\bm{\theta}^{i},\bm{\beta}\}\in\mathbb{R}^{61}$ follows the MANO [49] parameterization, where $\mathbf{R}_{h}^{i}$ is the rotation, $\mathbf{T}_{h}^{i}$ is the translation, $\bm{\theta}^{i}$ represents the pose and $\bm{\beta}$ represents shape parameters; $\mathbf{O}=\{\mathbf{o}^{i}\}_{i=1}^{l}$ is the object motion sequence, where each $\mathbf{o}^{i}=\{\mathbf{R}_{o}^{i},\mathbf{T}_{o}^{i}\}\in\mathbb{SE}(3)$ denotes the object’s pose, with $\mathbf{R}_{o}^{i}$ as the rotation and $\mathbf{T}_{o}^{i}$ as the translation. Additionally, we define the object geometry as a set of vertices $\mathbf{P}_{o}=\{\mathbf{p}^{n}_{o}\}_{n=1}^{N}\in\mathbb{R}^{N\times 3}$, where $N$ represents the number of object vertices. Given the input =$\mathcal{I}=\{\mathbf{G},\mathbf{h}^{1}_{l},\mathbf{h}^{1}_{r},\mathbf{o}^{1},\mathbf{P}_{o}\}$, our objective is to derive a subsequent hand-object interaction sequence $\langle\mathbf{H},\mathbf{O}\rangle$ that is not only consistent with the gaze but also exhibits natural and realistic interactions between the hand and the object.
IV-B GHO-Diffusion: Gaze-guided Hand-Object Diffusion
@@
The diffusion model contains a fixed forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ which gradually adds Gaussian noise to clean data $\mathbf{x}_{0}$ until it becomes pure noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and train a denoiser $p_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ to gradually denoise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to generate $\mathbf{x}_{0}$ in the reverse process. Instead of predicting $\bm{\epsilon}_{t}$ at each time step $t$, we follow MDM [50] to predict $\hat{\mathbf{x}}_{0}$ and optimizing $p_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ using the simple objective function:
@@
IV-B1 Stage-1: Object Dynamics Synthesis with Spatial-Temporal Gaze Encoding.
The gaze point locations are closely related to the object surface vertices, and since object motions have fewer degrees of freedom compared to hand motions, so we generate the object motions first. Gaze is a sparse and noisy signal, making it challenging to encode effectively. Drawing inspiration from GIMO [12], we design a gaze-guided spatial-temporal feature encoding module for object diffusion. We define $\mathbf{f}^{k}_{o}$ as the $k$-th local per-point object spatial feature extracted by PointNet++ [51]. Then according to the spatial relationship between gaze and object points, we can calculate gaze spatial feature:
@@
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are trainable weight matrices, and $d$ is the dimension of the feature.
We leverage the gaze spatial-temporal features $\mathbf{c}_{g}$, along with the object geometry $\mathbf{P}$ and initial object pose $\mathbf{o}^{1}$ as stage-1 diffusion condition and employ the objective function in Equation 4 as Stage-1 training loss to generate object motions $\hat{\mathbf{O}}$:
@@
where $\mathcal{L}_{\text{simple}}$ is defined by Equation 1.
The object translation loss $\mathcal{L}_{\text{trans}}$ is defined as:
$\mathcal{L}_{\text{trans}}=\sum\nolimits_{i=1}^{L}||\mathbf{T}_{o}^{i}-\hat{\mathbf{T}}_{o}^{i}||_{2},$ (5)
which enforces a constraint on the object’s position. The object vertices loss $\mathcal{L}_{\text{verts}}$ is defined as:
$\mathcal{L}_{\text{verts}}=\sum\nolimits_{i=1}^{l}||(\mathbf{R}_{o}^{i}(\mathbf{P})+\mathbf{T}_{o}^{i})-(\hat{\mathbf{R}}_{o}^{i}(\mathbf{P})+\hat{\mathbf{T}}_{o}^{i})||_{2},$ (6)
which enforces constraints on the object’s surface geometry, promoting finer details and ensuring more precise alignment. The smoothness loss $\mathcal{L}_{\text{smooth}}$ is defined as:
$\mathcal{L}_{\text{smooth}}=||\mathcal{H}^{1:l}-\mathcal{H}^{0:l-1}||_{2},$ (7)
which promotes the continuity and fluidity of the generated motion. By minimizing the difference between consecutive motion frames, this loss ensures that the generated object motion is smooth and avoids abrupt transitions between frames.
IV-B2 Stage-2: Hand Kinematic Synthesis with HOI Guidance